3x-4y=-9
-x-4y=-13

Answer by ptaylor(2198)

(1) 3x-4y=-9
(2) -x-4y=-13
We'll start by subtracting (2) from (1) and we get:
4x=4
x=1


Next, we'll substitute x=1 into (1) and we have:

3(1)-4y=-9
3-4y=-9 subtract 3 from both sides:
-4y=-12
y=3
CK
Substitute x=1 and y=3 into (1):
3(1)-4(3)=-9
-9=-9
Now substitute into (2)
-1-4(3)=-13
-13=-13

Hope this helps----ptaylor